OpenStudy (anonymous):

@Kristen17 An Unconditional acceptance into a graduate program at a university will be given to students whose GMAT scores plus 400 times the undergraduate grade point average is at least 1500. Robbins GMAT score was 600.What must her grade point average be in order to be unconditionally accepted into the program?

OpenStudy (anonymous):

Okay don't get me wrong...I might have this wrong but it kind of makes sense. 1500=400x+600 -600 -600 900=400x 900/400=400x/400 2.25=x

OpenStudy (thadyoung):

@Kristen17